OpenStudy (anonymous):

21/37=1/a +1/b +1/c a,b,c must be positive integers find a,b,c

OpenStudy (anonymous):

a=b=c=37/7

sam (.sam.):

\[\frac{1}{2}+\frac{1}{15}+\frac{1}{1110}\]

OpenStudy (anonymous):

how to solve?

OpenStudy (anonymous):

I guess it has infinite solutions....:S You have 1 ecuation with 3 unknowns :S

OpenStudy (anonymous):

but a,b,c must be positive integers

OpenStudy (anonymous):

but you didn't say that in your question..

OpenStudy (anonymous):

i'm sorry

OpenStudy (anonymous):

Then things change :D

sam (.sam.):

Its something to do with Egyptian fractions

OpenStudy (anonymous):

There is not single solution....you still have 3 unknowns and only one equation. Check this out; http://www.wolframalpha.com/input/?i=21%2F37%3D1%2Fa+%2B1%2Fb+%2B1%2Fc